### v1

# The numerical Amitsur group

Abstract.
The Amitsur subgroup of a variety with a group action measures the failure of the action to lift to the total spaces of its line bundles. We introduce the “numerical Amitsur group,” which is an approximation of the ordinary Amitsur subgroup that can be computed using only the Euler-Poincaré characteristic on the Picard group. As an application, we find a uniform upper bound on the exponent of the Amitsur subgroup that depends only on the dimension and arithmetic genus of the variety and is independent of the group. Finally, we compute Amitsur subgroups of toric varieties using these ideas.
Key words and phrases:
Amitsur subgroup, linearizations, equivariant birational geometry
2020 Mathematics Subject Classification:
14L30, 14E07
1. Introduction
Let $X$ be a smooth projective complex variety with an action of a finite group $G$. A line bundle $\mathcal{L}$ is linearizable if there exists a linear action of $G$ on the total space of $\mathcal{L}$ that is compatible with the action of $G$ on $X$. A choice of such action is a linearization of $\mathcal{L}$ and the set of isomorphism classes of line bundles together with a choice of linearization forms a group $\operatorname{Pic}(X,G)$.
Even if $g^{\ast}\mathcal{L}\cong\mathcal{L}$ for every $g\in G$, the action may not lift to the total space. There is an extension of $G$ by $\mathbb{C}^{\times}$ called the lifting group, which splits if and only if $\mathcal{L}$ is linearizable. We have an exact sequence
$0\to\operatorname{Hom}(G,\mathbb{C}^{\times})\to\operatorname{Pic}(X,G)\to\operatorname{Pic}(X)^{G}\xrightarrow{\partial}H^{2}(G,\mathbb{C}^{\times}).$
The image of $\partial$ is called the Amitsur subgroup and is denoted $\operatorname{Am}(X,G)$. Roughly speaking, the Amitsur subgroup measures the failure of line bundles to be $G$-linearizable. Alternatively, the Amitsur subgroup describes the set of all possible lifting groups of line bundles for $G$ acting on $X$.
Linearizations are central in Geometric Invariant Theory [MFK94] and the theta groups from the theory of abelian varieties are important examples of lifting groups [Mum08]. The Amitsur subgroup is an equivariant birational invariant and can be used to understand automorphism groups of Mori fiber spaces [BCDP23, Appendices]. An arithmetic version of the Amitsur subgroup sits inside the Brauer group of the base field and measures the failure of Galois descent for line bundles [Lie17]; indeed, this notion preceded the equivariant version we study here.
Let $J$ be a finite group acting on $\operatorname{Pic}(X)$ by group automorphisms. In this paper, the natural examples of such $J$ are subgroups of the automorphism group $\operatorname{Aut}(X)$, but many of our results apply more generally. Suppose that the Euler-Poincaré characteristic is $J$-invariant as a function of $\operatorname{Pic}(X)$; in other words $\chi(gD)=\chi(D)$ for every line bundle $D$ and element $g\in J$.
We introduce the numerical Amitsur group as the quotient
$\operatorname{Am}^{\chi}(X,J):=\operatorname{Pic}(X)^{J}/\operatorname{Pic}^{\chi}(X,J)$
of $\operatorname{Pic}(X)^{J}$ by the subgroup
$\operatorname{Pic}^{\chi}(X,J):=\left\langle\chi(D)\sum_{E\in J\cdot D}E\ \middle|\ D\in\operatorname{Pic}(X)\right\rangle,$
where $\chi(D)$ is the Euler-Poincaré characteristic of $D$, and $J\cdot D$ is the orbit of $D$ in $\operatorname{Pic}(X)$.
There are several reasons for introducing this notion. First of all, we show that the numerical Amitsur group is an “upper bound” for the ordinary Amitsur group:
Theorem 1.1 (cf. Theorem 5.2).
Let $G$ be a finite group and let $X$ be a smooth projective complex $G$-variety. There is a canonical surjection $\operatorname{Am}^{\chi}(X,G)\to\operatorname{Am}(X,G)$.
Secondly, the numerical Amitsur group is often easier to study. In practice, a group of automorphisms is often more complicated than the image of the action on the Picard group. This means that coarse estimates can be found for many groups at once. Moreover, if $\operatorname{Pic}(X)$ and $\chi$ are well understood, then the numerical Amitsur group can be effectively computed:
Theorem 1.2 (cf. Theorem 5.4).
If $\operatorname{Pic}(X)$ is finitely generated, then there is a finite subset $S$ such that
$\operatorname{Pic}^{\chi}(X,J):=\left\langle\chi(D)\sum_{E\in J\cdot D}E\ \middle|\ D\in S\right\rangle.$
To demonstrate these ideas in practice, we give several applications. We prove a uniform bound on the exponent of the Amitsur group using only the dimension and arithmetic genus:
Theorem 1.3 (cf. Theorem 4.2).
Let $G$ be a finite group and let $X$ be a smooth projective complex $G$-variety of dimension $n$. The exponent of $\operatorname{Am}(X,G)$ divides
$(1+(-1)^{n}p_{a})\operatorname{lcm}\{1,\ldots,n+1\}$
where $p_{a}$ is the arithmetic genus of $X$.
We also demonstrate how to compute the Amitsur group for toric varieties (Theorem 6.4). The numerical Amitsur group turns out to be a sharp upper bound in many cases:
Theorem 1.4 (cf. Proposition 6.6).
Let $X$ be a smooth projective complex Fano toric variety with a reductive automorphism group. Let $J$ be a subgroup of the image of $\operatorname{Aut}(X)$ in $\operatorname{Aut}(\operatorname{Pic}(X))$. There exists a finite group $G$ acting faithfully on $X$ and an isomorphism $\operatorname{Am}^{\chi}(X,J)\cong\operatorname{Am}(X,G)$.
The rest of the paper is structured as follows. In Section 2, we establish basic results about linearization and the Amitsur subgroup. In Section 3, we consider numerical polynomials and prove several technical results that will be used in subsequent sections. In Section 4, we explore how the Euler-Poincaré characteristic can be used to bound exponents of elements in the Amitsur group; in particular, we prove Theorem 1.3. In Section 5, we introduce the numerical Amitsur group and prove Theorems 1.1 and 1.2. In Section 6, we determine the Amitsur subgroups of toric varieties and prove Theorem 1.4. In Section 7, we work out several toric examples in detail; this serves both as a demonstration of the theory and to illustrate some of its subtleties.
2. Preliminaries
Let $G$ be a finite group. Let $X$ be a smooth projective complex $G$-variety. In other words, there is an action of $G$ on $X$ by morphisms of varieties. We do not assume the action is faithful.
2.1. Linearizations
We recall the notion of $G$-linearization of a vector bundle. See, for example, [Dol99, Bri18, MFK94].
Given vector bundles $\pi:E\to X$, $\pi^{\prime}:E^{\prime}\to X^{\prime}$, and a morphism $g:X\to X^{\prime}$ of varieties, an isomorphism of vector bundles lifting $g$ is an isomorphism $\varphi_{g}:E\to E^{\prime}$, which is linear on fibers, fitting into a Cartesian square
[…]
Equivalently, if $\mathcal{E}$ and $\mathcal{E}^{\prime}$ are the corresponding locally free sheaves, then $\varphi_{g}$ corresponds to an isomorphism $\mathcal{E}\cong g^{\ast}\mathcal{E}^{\prime}$ of locally free sheaves on $X$.
We will mainly be concerned with invertible sheaves and line bundles. There is a canonical homomorphism
$\alpha:\operatorname{Aut(X)}\to\operatorname{Aut}(\operatorname{Pic}(X))$
given by $g\mapsto(g^{\ast})^{-1}$. We use the notation $\operatorname{AutP}(X,G)$ to denote the image of the action of $G$ on $\operatorname{Pic}(X)$; in other words,
[…]
We write $\operatorname{AutP}(X):=\operatorname{AutP}(X,\operatorname{Aut}(X))$ for the case where $G=\operatorname{Aut}(X)$.
A locally free sheaf $\mathcal{E}$ is $G$-invariant if $g^{\ast}\mathcal{E}\cong\mathcal{E}$ for every $g\in G$. In the case where $\mathcal{L}$ is an invertible sheaf, $\mathcal{L}$ is $G$-invariant if and only if $[\mathcal{L}]\in\operatorname{Pic}(X)^{G}=\operatorname{Pic}(X)^{\operatorname{AutP}(X,G)}$.
Definition 2.1.
If $\mathcal{E}$ is a $G$-invariant locally free sheaf of finite rank then the lifting group $G_{\mathcal{E}}$ is the group of all isomorphisms of vector bundles (2.1) lifting $g$ over all $g\in G$.
The lifting group sits in a canonical exact sequence
(2.2) $1\to\operatorname{GL}_{r}(\mathbb{C})\to G_{\mathcal{E}}\to G\to 1,$
where $\operatorname{GL}_{r}(\mathbb{C})$ is the group of automorphisms of the locally free sheaf $\mathcal{E}$ of rank $r$ lifting the identity morphism of $X$.
In the case where $\mathcal{L}$ is an invertible sheaf, the canonical sequence is
(2.3) $1\to\mathbb{C}^{\times}\to G_{\mathcal{L}}\to G\to 1,$
where $\mathbb{C}^{\times}$ is a central subgroup corresponding to the automorphisms of the corresponding line bundle $L$ over the identity morphism of $X$.
Since the lifting group $G_{\mathcal{L}}$ acts on the line bundle $L$, each element $g$ in $G_{\mathcal{L}}$ induces a morphism $g_{\ast}=(g^{\ast})^{-1}:H^{i}(X,L)\to H^{i}(X,L)$ for every $i\geq 0$. Since $g_{\ast}\circ h_{\ast}=(g\circ h)_{\ast}$ for the pair $(X,L)$, this gives each $H^{i}(X,L)$ an action of $G_{\mathcal{L}}$. Indeed, we have the following:
Proposition 2.2.
The lifting group $G_{\mathcal{L}}$ acts on $H^{i}(X,\mathcal{L})$ linearly with $\mathbb{C}^{\times}$ in $G_{\mathcal{L}}$ identified with the action of the nonzero scalar matrices. If $G$ acts faithfully on $X$ and $\mathcal{L}$ is very ample, then the action of $G_{\mathcal{L}}$ on $H^{0}(X,\mathcal{L})$ is faithful.
For an invertible sheaf $\mathcal{L}$ in $\operatorname{Pic}(X)^{G}$ with $n+1=\dim H^{0}(X,\mathcal{L})\geq 2$, we obtain a rational map
$\varphi_{\mathcal{L}}:X\dasharrow\mathbb{P}^{n}.$
There is a canonical action of $G$ on $\mathbb{P}^{n}$ such that the rational map $\varphi_{\mathcal{L}}$ is $G$-equivariant. Moreover, if $\mathcal{L}$ is very ample, then the extension $G_{\mathcal{L}}$ is the pullback of
$1\to\mathbb{C}^{\times}\to\operatorname{GL}_{n+1}(\mathbb{C})\to\operatorname{PGL}_{n+1}(\mathbb{C})\to 1$
along the embedding $G\to\operatorname{PGL}_{n+1}(\mathbb{C})$.
Definition 2.3.
A $G$-invariant locally free sheaf $\mathcal{E}$ of finite rank is $G$-linearizable if the extension (2.2) defining $G_{\mathcal{E}}$ splits. A $G$-linearization of $\mathcal{E}$ is a choice of $G$-action on $\mathcal{E}$ that splits the sequence. The set of isomorphism classes of invertible sheaves with a choice of linearization form a group, which we denote $\operatorname{Pic}(X,G)$. The image of $\operatorname{Pic}(X,G)$ in $\operatorname{Pic}(X)$ will be denoted $\overline{\operatorname{Pic}}(X,G)$.
Proposition 2.4.
There is a canonical exact sequence
(2.4) $\begin{gathered}0\to\operatorname{Hom}(G,\mathbb{C}^{\times})\to\operatorname{Pic}(X,G)\to\operatorname{Pic}(X)^{G}\xrightarrow{\partial}H^{2}(G,\mathbb{C}^{\times}),\end{gathered}$
which is contravariantly functorial in both $X$ and $G$.
Proof.
This is well known. Perhaps, the most sophisticated way to see this is to use a Leray spectral sequence of stacks (see, e.g., [KT22, 3.1] and [PZ24, 1.3]). We will simply recall concrete interpretations of the maps leaving the remaining details to the reader.
Recall that the set of splittings of (2.3) is given by the group cohomology group $H^{1}(G,\mathbb{C}^{\times})$. Since the extension is central, this is canonically isomorphic to $\operatorname{Hom}(G,\mathbb{C}^{\times})$. The first morphism takes the set of splittings of $G_{\mathcal{O}_{X}}$ to the corresponding linearization.
The next morphism is the “forgetful” morphism which forgets the $G$-linearization of an invertible sheaf.
Recall that $H^{2}(G,\mathbb{C}^{\times})$ is canonically isomorphic to the group of extensions of $G$ by $\mathbb{C}^{\times}$. Therefore, for a $G$-invariant invertible sheaf $\mathcal{L}$, the element $\partial([\mathcal{L}])$ is simply the class of the extension (2.3) defining the lifting group $G_{\mathcal{L}}$. ∎
Since $H^{2}(G,\mathbb{C}^{\times})$ is finite and torsion, it is immediately clear that for every $G$-invariant invertible sheaf $\mathcal{L}$, there exists some positive integer $d$ such that $d\partial(\mathcal{L})=0$.
Definition 2.5.
Suppose $\mathcal{L}$ is a $G$-invariant invertible sheaf on $X$. The Amitsur period of $\mathcal{L}$, denoted $\mathsf{m}(X,G;\mathcal{L})$, is the exponent of the element $\partial(\mathcal{L})$ in the group $H^{2}(G,\mathbb{C}^{\times})$. We use the shorthand $\mathsf{m}(\mathcal{L})$ when the $G$-variety is clear. The Amitsur period of the $G$-variety $X$, denoted $\mathsf{m}(X,G)$, is the least common multiple of $\mathsf{m}(X,G;\mathcal{L})$ over all line bundles $[\mathcal{L}]\in\operatorname{Pic}(X)^{G}$.
An arithmetic version of the following proposition, in the case of $H^{0}$, can be found in [CS21, Proposition 7.1.15(i)].
Proposition 2.6.
If $[\mathcal{L}]$ is in $\operatorname{Pic}(X)^{G}$ and there exists a $G_{\mathcal{L}}$-subrepresentation $V$ of $H^{i}(X,\mathcal{L})$ of dimension $d$, then $d\partial(\mathcal{L})=0$. In other words, $\mathsf{m}(\mathcal{L})$ divides $d$.
Proof.
Let $G_{d\mathcal{L}}$ be the group in the extension $d\partial(\mathcal{L})=\partial(\mathcal{L}^{\otimes d})$ in $H^{2}(G,\mathbb{C}^{\times})$. In other words, we have the following commutative diagram of exact sequences:
[…]
where the morphism $\mathbb{C}^{\times}\to\mathbb{C}^{\times}$ is the power map $x\mapsto x^{d}$. We want to show the bottom sequence splits.
We have a $d$-dimensional representation $V$ of $G_{\mathcal{L}}$ whose restriction to $\mathbb{C}^{\times}$ corresponds to multiplication by scalar matrices. The representation $\Lambda^{d}V$ is a $1$-dimensional representation of $G_{\mathcal{L}}$ whose restriction to $\mathbb{C}^{\times}$ is the power map $x\mapsto x^{d}$. Therefore, we have a homomorphism $G_{\mathcal{L}}\to\mathbb{C}^{\times}$ which factors through $G_{d\mathcal{L}}$ in the diagram above. The resulting quotient morphism $G_{d\mathcal{L}}\to\mathbb{C}^{\times}$ gives the desired splitting. ∎
The following fact, apparently first observed by A. Kuznetsov, is the foundation for this paper. An arithmetic version of this can be found in [CS21, Proposition 7.1.15(ii)].
Proposition 2.7.
Let $\chi(\mathcal{L})$ be the Euler-Poincaré characteristic of $\mathcal{L}$. We have $\chi(\mathcal{L})\partial(\mathcal{L})=0$. In other words, $\mathsf{m}(\mathcal{L})$ divides $\chi(\mathcal{L})$.
Proof.
In view of Proposition 2.6, we have $\left(\dim H^{i}(X,\mathcal{L})\right)\partial(\mathcal{L})=0$ for every $i$. Since $\chi(\mathcal{L})$ is an alternating sum of $\dim H^{i}(X,\mathcal{L})$, the result follows immediately. ∎
From [BCDP23, Proposition 2.11], we have the following.
Proposition 2.8.
The canonical line bundle has a canonical linearization.
2.2. Amitsur subgroup
Our main interest is the Amitsur subgroup
$\operatorname{Am}(X,G):=$ $\operatorname{im}\left(\partial:\operatorname{Pic}(X)^{G}\to H^{2}(G,\mathbb{C}^{\times})\right)$
$\cong$ $\operatorname{coker}\left(\operatorname{Pic}(X,G)\to\operatorname{Pic}(X)^{G}\right)$
first named in [BCDP23]. This is an equivariant analog of the arithmetic version first named in [Lie17], although its study goes back decades.
In this paper, we are less interested in the specific embedding into $H^{2}(G,\mathbb{C}^{\times})$ and are more interested in its description as a quotient of $\operatorname{Pic}(X)^{G}$. Therefore, we will more frequently refer to it as the Amitsur group rather than the Amitsur subgroup.
The arithmetic interest is related to the fact that it a birational invariant. This is true in the equivariant context as well by [BCDP23, Theorem A.1]:
Theorem 2.9.
If $G$ acts faithfully on $X$, then $\operatorname{Am}(X,G)$ is a $G$-equivariant birational invariant of smooth projective $G$-varieties.
The following Theorem is due to Dolgachev [Dol99]:
Theorem 2.10.
If $X$ is a curve and $G$ acts faithfully, then $\operatorname{Am}(X,G)=H^{2}(G,\mathbb{C}^{\times})$.
Suppose $G$ and $H$ are finite groups and $\varphi:G\to H$ is a group homomorphism. Suppose $X$ is a $G$-variety, $Y$ is an $H$-variety and $f:X\to Y$ is a $\varphi$-equivariant morphism. In other words, $f(g\cdot x)=\varphi(g)f(x)$ for all $g\in G$. Then there is an induced map
$f^{\ast}:\operatorname{Am}(Y,H)\to\operatorname{Am}(X,G)$
using the functoriality of the exact sequence (2.4). Indeed, $\operatorname{Am}(-,-)$ is a contravariant functor from the category of varieties with a finite group action to the category of abelian groups.
Proposition 2.11.
If $f:X\to Y$ is a $G$-equivariant morphism, then the induced morphism $f^{\ast}:\operatorname{Am}(Y,G)\to\operatorname{Am}(X,G)$ is injective. If $f$ has a $G$-equivariant section, then $\operatorname{Am}(Y,G)$ is a direct summand.
Proof.
See [BCDP23, Lemma A.6], where faithfulness is assumed in the statement, but not needed in the proof. ∎
Even restriction to a subgroup $H\subset G$ can be a subtle operation. Indeed, in Remark 7.3 below, we see that it is possible for $\operatorname{Am}(X,G)$ to be trivial, while $\operatorname{Am}(X,H)$ is nontrivial. However, we have the following. Recall that $\operatorname{AutP}(X,G)$ is the image of the $G$-action on $\operatorname{Pic}(X)$.
Proposition 2.12.
Suppose that $\varphi:H\to G$ is a group homomorphism and $\operatorname{AutP}(X,G)=\operatorname{AutP}(X,H)$. Then the induced homomorphism $\operatorname{Am}(X,G)\to\operatorname{Am}(X,H)$ is surjective.
Proof.
This follows by the description of $\operatorname{Am}(X,G)$ as the cokernel of the map
$\operatorname{Pic}(X,G)\to\operatorname{Pic}(X)^{G}.$
By our hypothesis, $\operatorname{Pic}(X)^{G}=\operatorname{Pic}(X)^{H}$. A line bundle is $H$-linearizable if it is $G$-linearizable, so the image of $\operatorname{Pic}(X,G)$ is contained in the image of $\operatorname{Pic}(X,H)$. ∎
Proposition 2.13.
If $\varphi:H\to G$ is a group homomorphism then the kernel of the homomorphism $\operatorname{Am}(X,G)\to\operatorname{Am}(X,H)$ is the set of classes $[\mathcal{L}]$ where $\mathcal{L}$ is a $G$-invariant line bundle on $X$ such that $\varphi$ factors through $G_{\mathcal{L}}\to G$.
Proof.
The kernel of the morphism $H^{2}(G,\mathbb{C}^{\times})\to H^{2}(H,\mathbb{C}^{\times})$ is the set of classes that are trivialized when pulled back along $\varphi$. ∎
3. Numerical Polynomials
We recall a special case of Snapper’s Theorem [Kle66]. Given a set of line bundles $\mathcal{L}_{1},\ldots,\mathcal{L}_{r}$ on a variety $X$, the Euler-Poincaré characteristic
$(a_{1},\ldots,a_{r})\mapsto\chi\left(\mathcal{L}_{1}^{\otimes a_{1}}\otimes\cdots\otimes\mathcal{L}_{r}^{\otimes a_{r}}\right)$
is a numerical polynomial. In this section, we recall some well known facts about numerical polynomials as well as prove some new facts which will be important below.
We recall that a numerical polynomial or integer-valued polynomial is a polynomial $f\in\mathbb{Q}[x_{1},\ldots,x_{r}]$ such that $f(x_{1},\ldots,x_{r})$ is an integer for all tuples $(x_{1},\ldots,x_{r})\in\mathbb{Z}^{r}$. In this section, we write $\mathbf{x}=(x_{1},\ldots,x_{r})$ for elements of $\mathbb{Z}^{r}$.
Recall that the binomial coefficient
$\binom{x}{a}=\frac{x(x-1)\cdots(x-a+1)}{a!}$
is a numerical polynomial in $\mathbb{Q}[x]$ of degree $a$. For $\mathbf{x}\in\mathbb{Z}^{r}$ and $\mathbf{a}\in\mathbb{N}^{r}$ we write
$\binom{\mathbf{x}}{\mathbf{a}}:=\binom{x_{1}}{a_{1}}\cdots\binom{x_{r}}{a_{r}}$
where each $\binom{x_{i}}{a_{i}}$ is a binomial coefficient.
If $f$ is a numerical polynomial of degree $n$ in $r$ variables, then there exists a unique expression
$f(\mathbf{x})=\sum_{\mathbf{a}}c_{\mathbf{a}}\binom{\mathbf{x}}{\mathbf{a}}$
where the sum runs over all $\mathbf{a}$ satisfying the relations
[…]
and $c_{\mathbf{a}}$ are integers depending on $f$.
For each index $i$, we have the forward difference operator
$(\Delta_{i}f)(\mathbf{x})=f(\mathbf{x}+\mathbf{e}_{i})-f(\mathbf{x})$
where $\mathbf{e}_{i}$ is the $i$th standard basis vector in $\mathbb{Z}^{r}$. These operators can be iterated to obtain
$(\Delta_{i}^{a}f)(\mathbf{x})=\sum_{k=0}^{a}(-1)^{a-k}\binom{a}{k}f(\mathbf{x}+k\mathbf{e}_{i})$
For $\mathbf{a}\in\mathbb{N}^{r}$, we define the operator
$\Delta^{\mathbf{a}}f:=\Delta_{1}^{a_{1}}\cdots\Delta_{r}^{a_{r}}f.$
We recall the multivariate Newton-Gregory interpolation formula:
Proposition 3.1.
If $f$ is an integer-valued polynomial in $r$ variables of degree $n$, then
$f(\mathbf{y}+\mathbf{x})=\sum_{\mathbf{a}}\binom{\mathbf{x}}{\mathbf{a}}(\Delta^{\mathbf{a}}f)(\mathbf{y})$
where the sum is over $\mathbf{a}\in\mathbb{N}^{r}$ such that $\sum_{i=1}^{r}a_{i}\leq n$.
The following is the key new concept we need for producing finite generating sets for the numerical Amitsur group.
Definition 3.2.
A subset $S\subseteq\mathbb{Z}^{r}$ is integrally-poised for polynomials of degree $n$ if there exists a set of numerical polynomials $\{c_{\mathbf{a}}\}_{\mathbf{a}\in S}$ such that
$f(\mathbf{x})=\sum_{\mathbf{a}\in S}c_{\mathbf{a}}(\mathbf{x})f(\mathbf{a})$
for all polynomials $f$ of degree $n$ and all $\mathbf{x}\in\mathbb{Z}^{r}$.
Our main example of an integrally-poised subset is the lattice simplex:
Proposition 3.3.
The set
$S_{n}=\left\{\mathbf{a}\in\mathbb{Z}^{r}\ \middle|\ a_{1}\geq 0,\ \ldots,\ a_{r}\geq 0,\ \sum_{i=1}^{r}a_{i}\leq n\right\}$
is integrally-poised for polynomials of degree $n$.
Proof.
Suppose $f(\mathbf{x})$ is a multivariate integer-valued polynomial in $r$ variables of degree $n$. From above, we see that $(\Delta^{\mathbf{a}}f)(0)$ is a specific integral linear combination of $f(\mathbf{b})$ where $0\leq b_{i}\leq a_{i}$ for every index $i$. From the Newton-Gregory interpolation formula, we see that
$f(\mathbf{x})=\sum_{\mathbf{a}\in S_{n}}c_{\mathbf{a}}(\mathbf{x})f(\mathbf{a})$
where each $c_{\mathbf{a}}$ is a multivariate integer-valued polynomial that only depends on the number of variables $r$ and degree $n$. ∎
Finally, we obtain the result needed for our application to Amitsur groups.
Proposition 3.4.
Suppose $S\subseteq\mathbb{Z}^{r}$ is integrally poised for polynomials of degree $n+1$. If $g$ is an integer-valued polynomial degree $n$, then we have equality of the subgroups
(3.1) $\langle g(\mathbf{x})\mathbf{x}\mid\mathbf{x}\in\mathbb{Z}^{n}\rangle=\langle g(\mathbf{x})\mathbf{x}\mid\mathbf{x}\in S\rangle$
of $\mathbb{Z}^{r}$.
Proof.
By assumption, we have polynomials $c_{\mathbf{a}}(\mathbf{x})$ such that
$f(\mathbf{x})=\sum_{\mathbf{a}\in S}c_{\mathbf{a}}(\mathbf{x})f(\mathbf{a})$
for every polynomial $f$ of degree $n+1$. Applying the formula to $f(\mathbf{x})=g(\mathbf{x})x_{i}$ for each $i=1,\ldots,r$ we recover the formula
$g(\mathbf{x})\mathbf{x}=\sum_{\mathbf{a}\in S}c_{\mathbf{a}}(\mathbf{x})g(\mathbf{a})\mathbf{a}.$
This shows that every generator on the left hand side of (3.1) is contained in the right hand side. ∎
3.1. Univariate Polynomials
Let $f(x)$ be a univariate numerical polynomial of degree $n$. Recall that $f$ has a unique expression
$f(x)=\sum_{i=0}^{n}a_{i}\binom{x}{i}$
for integers $a_{0},\ldots,a_{n}$. By convention, we set $a_{i}=0$ for all $i<0$ and $i>n$.
Lemma 3.5.
$\gcd\{f(k)\}:=\gcd\{f(k)\mid k\in\mathbb{Z}\}=\gcd\{a_{0},\ldots,a_{n}\}$
Proof.
From Prop 3.3, we see that $\gcd\{f(k)\mid k\in\mathbb{Z}\}$ is equal to the expression $\gcd\{f(0),f(1),\ldots,f(n)\}$. Consider the equations
(3.2) $f(j)=\sum_{i=0}^{n}a_{i}\binom{j}{i}$
for integers $0\leq j\leq n$. This a binomial transform, thus each $a_{i}$ can be written as an integral linear combination of $f(0),\ldots,f(n)$. In particular, the two sets have the same $\gcd$.
For those not familiar with binomial transforms, we sketch an argument. We can view 3.2 as a matrix equation where we’ve multiplied a vector with entries $a_{i}$ by a matrix with entries $\binom{j}{i}$ to obtain a vector with entries $f(j)$. Since $\binom{j}{i}=0$ for all $j<i$ and $\binom{i}{i}=1$ for all $i$, the matrix is invertible over the integers. ∎
Lemma 3.6.
$kf(k)=\left(\sum_{i=0}^{n}i(a_{i}+a_{i-1})\binom{k}{i}\right)+(n+1)a_{n}\binom{k}{n+1}$
Proof.
Rearranging the identity
$\binom{k}{i+1}=\frac{k-i}{i+1}\binom{k}{i}$
we obtain
$k\binom{k}{i}=(i+1)\binom{k}{i+1}+i\binom{k}{i}.$
The result follows immediately. ∎
Applying Lemma 3.5 to Lemma 3.6, we obtain:
Corollary 3.7.
An integer $m$ divides $\gcd\{kf(k)\mid k\in\mathbb{Z}\}$ if and only if the integers $a_{0},\ldots,a_{n}$ satisfy the congruences
$a_{1}+a_{0}$ $\equiv 0\mod{m}$
$2(a_{2}+a_{1})$ $\equiv 0\mod{m}$
$3(a_{3}+a_{2})$ $\equiv 0\mod{m}$
$\vdots$
$k(a_{k}+a_{k-1})$ $\equiv 0\mod{m}$
$\vdots$
$n(a_{n}+a_{n-1})$ $\equiv 0\mod{m}$
$(n+1)a_{n}$ $\equiv 0\mod{m}.$
Lemma 3.8.
$\frac{\gcd\{kf(k)\}}{\gcd\{f(k)\}}$ divides $\operatorname{lcm}\{1,\ldots,n+1\}$.
Proof.
It follows immediately from the definitions that $\gcd\{f(k)\}$ divides $\gcd\{kf(k)\}$. Now we apply Corollary 3.7. Since the equations are homogeneous, we may safely divide by $\gcd\{f(k)\}$ and assume that $a_{0},\ldots,a_{n}$ have no common factor. Let $p$ be a prime and $s$ be a positive integer such that $p^{s}$ divides $\gcd\{kf(k)\}$. It remains to prove that $p^{s}\leq n+1$.
Suppose otherwise; that $p^{s}>n+1$. Consider $10<k\leq n+$ 1. Write $k=p^{r}m$ where $r$ is a non-negative integer and $m$ is a positive integer coprime to $p$. Since $k<p^{s}$, we have $r<s$. Therefore, we have the following:
$ka_{k}$ $\equiv-ka_{k-1}\pmod{p^{s}}$
$p^{r}ma_{k}$ $\equiv-p^{r}ma_{k-1}\pmod{p^{s}}$
$ma_{k}$ $\equiv-ma_{k-1}\pmod{p^{s-r}}$
$ma_{k}$ $\equiv-ma_{k-1}\pmod{p}$
$a_{k}$ $\equiv-a_{k-1}\pmod{p}.$
From $k=n+1$, we conclude that $a_{n}$ is divisible by $p$ since $a_{n+1}=0$. From $k=n$, we then conclude that $a_{n-1}$ is also divisible by $p$. Continuing in this way, we conclude that $a_{0},\ldots,a_{n}$ are all divisible by $p$. This contradicts that they have no common factor. ∎
Remark 3.9.
The expression
$\operatorname{lcm}\{1,\ldots,n+1\}$
is exactly the exponent of the symmetric group $S_{n+1}$. This function has many other interpretations [OEI25, A003418]. In particular, the logarithm $\psi(x)=\log\operatorname{lcm}\{1,\ldots,x\}$ is the second Chebyshev function. One can show that $\psi(x)\sim x$ and certain more refined estimates turn out to be equivalent to the Riemann Hypothesis [Dav00, §18]. (We thank F. Thorne for pointing out this reference.)
4. Upper bounds on Amitsur periods
Suppose $X$ is a smooth projective variety with an action of a finite group $G$. Recall that, if $\mathcal{L}$ is $G$-invariant line bundle, then the Amitsur period $\mathsf{m}(X,G;\mathcal{L})=\mathsf{m}(\mathcal{L})$ is the exponent of $\partial[\mathcal{L}]$ in the Amitsur group $\operatorname{Am}(X,G)$. The value $\mathsf{m}(X,G)$ is the exponent of the entire group $\operatorname{Am}(X,G)$.
Lemma 4.1.
If $X$ is a smooth $G$-variety and $\mathcal{L}$ is a $G$-invariant line bundle class, then $\mathsf{m}(\mathcal{L})$ divides $\gcd\{kf(k)\}$ where $f(k)=\chi(\mathcal{L}^{\otimes{k}})$.
Proof.
By Proposition 2.7, we have $\chi(\mathcal{L})\partial(\mathcal{L})=0$ for every $G$-invariant line bundle $\mathcal{L}$. Using additive notation with a divisor $D$ representing the class of $\mathcal{L}$, this means that
$\left\langle\chi(kD)k[D]\ \middle|\ k\in\mathbb{Z}\right\rangle$
is a subgroup of the image $\overline{\operatorname{Pic}}(X,G)$ of $\operatorname{Pic}(X,G)$ in $\operatorname{Pic}(X)^{G}$. This subgroup is generated by $\gcd\{kf(k)\}[D]$. ∎
4.1. A uniform bound
We are now in a position to prove the uniform bound stated in the introduction.
Theorem 4.2.
Let $X$ be a smooth variety of dimension $n$ with arithmetic genus $p_{a}$. For any finite group $G$, the Amitsur period $\mathsf{m}(X,G)$ divides
$(1+(-1)^{n}p_{a})\operatorname{lcm}\{1,\ldots,n+1\}.$
Proof.
The choice of finite group is irrelevant. We consider the polynomial $f(k)=\chi(kD)$ where $D$ represents a $G$-invariant divisor class. Observe that $f(0)=\chi(\mathcal{O}_{X})=(1+(-1)^{n}p_{a})$. Therefore $\gcd\{f(k)\}$ divides $f(0)$. The result follows immediately from Lemmas 3.8 and 4.1. ∎
Remark 4.3.
For any positive integer $m$, there exists a choice of finite group $G$ such that $\operatorname{Am}(\mathbb{P}^{m-1},G)\cong\mathbb{Z}/m\mathbb{Z}$. If $n\geq m-1$, then we may construct $X\cong\mathbb{P}^{m-1}\times\mathbb{P}^{n-m+1}$ where $G$ acts trivially on $\mathbb{P}^{n-m+1}$. We see that $X$ has dimension $n$ and $\operatorname{Am}(X,G)\cong\mathbb{Z}/m\mathbb{Z}$ has exponent $m$. This construction works for any $m\leq n+1$; in particular for any prime power $\leq n+1$. Therefore, for varieties of genus 00, the bound from Theorem 4.2 is multiplicatively sharp.
Remark 4.4.
The bound is vacuous for curves of genus $1$ since $\chi(\mathcal{O}_{X})=0$. This is not surprising since there is no uniform bound for any such curve if one does not fix the group $G$. Indeed, for every elliptic curve and every prime $p$, there is a group of automorphisms $G$ isomorphic to $C_{p}^{2}$ that acts by translations. Recall that $H^{2}(G,\mathbb{C}^{\times})\cong\mathbb{Z}/p\mathbb{Z}$. By Theorem 2.10, we see that $\operatorname{Am}(X,G)=\mathbb{Z}/p\mathbb{Z}$. Therefore, the period $\mathsf{m}(X,G)$ can be made arbitrarily large by appropriate choice of $G$.
Remark 4.5.
The bound is not necessarily sharp for higher genus curves. By Theorem 2.10, if $G$ acts faithfully then $\mathsf{m}(X,G)$ is the exponent of $H^{2}(G,\mathbb{C}^{\times})$. The Hurwitz bound states that $|G|\leq 84(p_{a}-1)$ for a group acting faithfully on a curve $X$. However, if the $p$-primary component of $H^{2}(G,\mathbb{C}^{\times})$ is non-zero, then $G$ must have order at least $p^{2}$.
4.2. Bounds on the Amitsur period for specific divisors
The remainder of this section demonstrate how one can use the numerical results from the previous section to bound the Amitsur periods of specific divisors using intersection-theoretic information. In order to do this, we use expressions
(4.1) $f(k)=\chi(kD)=\sum_{i=0}^{n}a_{i}\binom{k}{i}$
where $a_{0},\ldots,a_{n}$ are constants depending on $D$.
Proposition 4.6.
Suppose $X$ is a curve with genus $p_{a}$ and $D$ is a divisor whose class is in $\operatorname{Pic}(X)^{G}$.
  1. (a)
    $\mathsf{m}(D)$ divides $\gcd\{kf(k)\}$.
  2. (b)
    $\gcd\{f(k)\}=\gcd\{(1-p_{a}),\deg(D)\}$
  3. (c)
    $\gcd\{kf(k)\}=\gcd\{\deg(D)+(1-p_{a}),2\deg(D)\}$
  4. (d)
    $\frac{\gcd\{kf(k)\}}{\gcd\{f(k)\}}=\begin{cases}2&\textrm{$1-p_{a}$ and $\deg(D)$ have same parity}\\ 1&\textrm{otherwise}.\end{cases}$.
Proof.
The Riemann-Roch theorem gives us
$\chi(kD)=\deg(kD)+(1-p_{a})=\left(\deg(D)\right)k+(1-p_{a}).$
Thus, $a_{0}=(1-p_{a})$ and $a_{1}=\deg(D)$ in (4.1) and the results follow from the results in the previous section. ∎
Proposition 4.7.
Suppose $X$ is a surface with arithmetic genus 00 and $D$ is a divisor whose class is in $\operatorname{Pic}(X)^{G}$.
  1. (a)
    $\mathsf{m}(D)$ divides $6$.
  2. (b)
    $\mathsf{m}(D)$ divides $\chi(D)$.
  3. (c)
    If $2$ divides $\mathsf{m}(D)$, then $D^{2}$ and $D\cdot K_{X}$ are both even, but are distinct modulo $4$.
  4. (d)
    If $3$ divides $\mathsf{m}(D)$, then $D\cdot K_{X}\equiv 0\mod{3}$ and $D^{2}\equiv 1\mod{3}$.
Proof.
Since $\chi(0)=1+p_{a}=1$, we conclude $\mathsf{m}(D)$ divides $6$ from Theorem 4.2. The second bound is Proposition 2.7, but we include it here for completeness. For the remaining bounds, we use the fact that $\mathsf{m}(D)$ divides $\gcd\{kf(k)\}$ by Lemma 4.1.
The usual Riemann-Roch theorem for surfaces is
$\chi(D)=1+p_{a}+\frac{1}{2}\left(D^{2}-D\cdot K_{X}\right)$
From this, we determine the coefficients
$a_{0}$ $=\chi(0)=1+p_{a}=1$
$a_{1}$ $=\frac{1}{2}\left(D^{2}-D\cdot K_{X}\right)$
$a_{2}$ $=D^{2}$
for the binomial coefficients in (4.1). Thus, by Corollary 3.7, the multiplicity $\mathsf{m}(D)$ divides all of the following expressions:
(4.2) $a_{1}+a_{0}$ $=\frac{1}{2}\left(D^{2}-D\cdot K_{X}\right)+1$
(4.3) $2(a_{2}+a_{1})$ $=3D^{2}-D\cdot K_{X}$
(4.4) $3a_{2}$ $=3D^{2}$
Suppose $\gcd\{kf(k)\}$ is divisible by $2$. Immediately, we see that $D^{2}$ must be even. Using expression (4.3), we conclude that $D\cdot K_{X}$ is also even. Next, expression (4.2) implies that $\frac{1}{2}\left(D^{2}-D\cdot K_{X}\right)$ is odd.
Suppose $\gcd\{kf(k)\}$ is divisible by $3$. From expression (4.3), we see that $D\cdot K_{X}\equiv 0\mod{3}$. Now, from expression (4.2), we have $D^{2}+2\equiv 0\mod{3}$. ∎
Proposition 4.8.
Suppose $X$ is a threefold with arithmetic genus 00 and $D$ is a divisor whose class is in $\operatorname{Pic}(X)^{G}$.
  1. (a)
    $\mathsf{m}(D)$ divides $12$.
  2. (b)
    $\mathsf{m}(D)$ divides $\chi(D)$.
  3. (c)
    $2D^{3}\equiv 6\pmod{m}$ and $D^{2}\cdot K_{X}\equiv 4\pmod{2m}$
Proof.
As before, the first two bounds follow from Theorem 4.2 and Proposition 2.7. Using Hirzebruch-Riemann-Roch for threefolds, we find
$\chi(kD)=\frac{1}{6}(kD)^{3}-\frac{1}{4}(kD)^{2}\cdot K_{X}+\frac{1}{12}(kD)\cdot\left(K_{X}^{2}+c_{2}\right)+1.$
Viewing this as a polynomial in $k$ in the binomial basis, we obtain the following coefficients
$a_{0}$ $=1$
$a_{1}$ $=\frac{1}{6}D^{3}-\frac{1}{4}D^{2}\cdot K_{X}+\frac{1}{12}D\cdot\left(K_{X}^{2}+c_{2}\right)$
$a_{2}$ $=D^{3}-\frac{1}{2}D^{2}\cdot K_{X}$
$a_{3}$ $=D^{3}$
as in (4.1). Thus, $\mathsf{m}(D)$ divides all of the following expressions:
(4.5) $a_{1}+a_{0}$ $=\frac{1}{6}D^{3}-\frac{1}{4}D^{2}\cdot K_{X}+\frac{1}{12}D\cdot\left(K_{X}^{2}+c_{2}\right)+1$
(4.6) $2(a_{2}+a_{1})$ $=\frac{7}{3}D^{3}-\frac{3}{2}D^{2}\cdot K_{X}+\frac{1}{6}D\cdot\left(K_{X}^{2}+c_{2}\right)$
(4.7) $3(a_{3}+a_{2})$ $=6D^{3}-\frac{3}{2}D^{2}\cdot K_{X}$
(4.8) $4a_{3}$ $=4D^{3}.$
Let $m$ be a divisor of $\mathsf{m}(D)$. Subtracting twice (4.5) from (4.6), and multiplying (4.7) by $2$, we obtain the congruences
(4.9) $\chi(D)$ $\equiv 0\mod{m}$
(4.10) $2\displaystyle 2D^{3}-D^{2}\cdot K_{X}-$ 2 $\equiv 0\mod{m}$
(4.11) $12D^{3}-3D^{2}\cdot K_{X}$ $\equiv 0\mod{2m}$
(4.12) $4D^{3}$ $\equiv 0\mod{m}.$
Taking twice the sum of (4.10) and (4.12), then subtracting (4.11), we find
$D^{2}\cdot K_{X}\equiv 4\mod{2m}.$
Substituting this back into the congruences above, we find that
$4D^{3}\equiv 0\pmod{m}\textrm{ and }2D^{3}\equiv 6\pmod{m}.$
Since $m$ divides $12$, the first congruence is redundant. ∎
5. Numerical Amitsur group
Suppose $X$ is a smooth projective variety. Let $G$ be a finite group acting on $\operatorname{Pic}(X)$ by group automorphisms such that $\chi(D)=\chi(g^{\ast}D)$ for all $g\in G$ and $[D]\in\operatorname{Pic}(X)$. The reader should imagine that $G$ is $\operatorname{AutP}(X,G)$ from above, although we do not necessarily assume that $G$ comes from an action on $X$.
We use the following shorthand for orbit sums. For $[D]\in\operatorname{Pic}(X)$, write
$\Sigma_{G}(D):=\sum_{[E]\in G\cdot[D]}[E]$
where $G\cdot[D]$ is the $G$-orbit of $[D]$ in $\operatorname{Pic}(X)$. Alternatively,
$\Sigma_{G}(D)=\sum_{g\in G/H}g[D]$
where $H$ is the stabilizer of $[D]$ in $G$. Observe that $\Sigma_{G}(D)$ is always in $\operatorname{Pic}(X)^{G}$ even though $[D]$ may not be.
Definition 5.1.
The numerical Amitsur group $\operatorname{Am}^{\chi}(X,G)$ is the cokernel in the short exact sequence
(5.1) $0\to\operatorname{Pic}^{\chi}(X,G)\to\operatorname{Pic}(X)^{G}\to\operatorname{Am}^{\chi}(X,G)\to 0$
where $\operatorname{Pic}^{\chi}(X,G)$ is the subgroup
(5.2) $\operatorname{Pic}^{\chi}(X,G):=\left\langle\chi(D)\Sigma_{G}(D)\ \middle|\ [D]\in\operatorname{Pic}(X)\right\rangle.$
A divisor class $[D]$ is numerically $G$-split if and only if $[D]\in\operatorname{Pic}^{\chi}(X,G)$. (We anticipate that $\operatorname{Pic}^{\chi}(X,G)$ will also be useful in the arithmetic setting, which explains the terminology “split” instead of “linearizable.”)
The notation $\operatorname{Pic}^{\chi}(X,G)$ suggests that it is a “numerical” version of $\operatorname{Pic}(X,G)$, but the latter is not a subgroup of $\operatorname{Pic}(X)$. Instead, $\operatorname{Pic}^{\chi}(X,G)$ is more analogous to the group $\overline{\operatorname{Pic}}(X,G)$ of isomorphism classes of linearizable line bundles where the particular choice of linearization is not part of the data.
We now show, as stated in the introduction, that the numerical Amitsur group is an “upper bound” for the ordinary Amitsur group.
Theorem 5.2.
There is a canonical surjection
$\operatorname{Am}^{\chi}(X,\operatorname{AutP}(X,G))\to\operatorname{Am}(X,G)$
for every finite group $G$ acting on $X$.
Proof.
Using (2.4), it suffices to show that every element of $\operatorname{Pic}^{\chi}(X,G)$ is linearizable. Suppose $[D]$ is in $\operatorname{Pic}^{\chi}(X,G)$. Since tensor products of linearizable invertible sheaves are linearizable, it suffices to assume that
$[D]=\chi(E)\sum_{[F]\in G\cdot[E]}[F]$
for some divisor class $[E]\in\operatorname{Pic}(X)$. Let $H$ be the stabilizer of $[E]$ in $\operatorname{Pic}(X)$. Then $\chi(E)[E]$ is $H$-linearizable by Proposition 2.6. Thus $[D]$ is $G$-linearizable by Lemma 5.3 below. ∎
Lemma 5.3.
Suppose $H$ is the stabilizer in $G$ of $[D]\in\operatorname{Pic(X)}$. If $D$ is $H$-linearizable, then $\Sigma_{G}(D)$ is $G$-linearizable.
Proof.
Let $\mathcal{L}$ be the corresponding invertible sheaf. Let $t_{1},\ldots,t_{s}$ be a system of distinct representatives for the left cosets in $G/H$ and define
$\mathcal{F}:=\bigoplus_{i=1}^{s}t_{i}^{\ast}\mathcal{L}.$
Observe that the determinant of $\mathcal{F}$ has the same class as $[D]$. If $\mathcal{F}$ has a linearization, then so must its determinant. Observe that the determinant is $\Sigma_{G}(D)$. Thus, it remains to exhibit a $G$-linearization on $\mathcal{F}$. We do so using a construction inspired from the “induced representation” from ordinary representation theory.
We change to the language of total spaces. Let $L$ (resp. $F$) be the total space of $\mathcal{L}$ (resp. $\mathcal{F}$) and let $\pi:L\to X$ be the projection. Let $L^{s}$ be the $s$-fold product of $L$ (as a variety) and let $p_{i}:L^{s}\to L$ denote the $i$th projection. For each $g\in G$ and index $i$ there is a unique index $g(i)$ and element $h_{g,i}\in H$ such that $gt_{i}=t_{g(i)}h_{g,i}$. We define an action of $G$ on $L^{s}$ by requiring that
$p_{i}(g\cdot\ell):=h_{g,g^{-1}(i)}p_{g^{-1}(i)}(\ell)$
for each index $i$. One checks that this gives a well-defined $G$-action.
Observe that $F$ is the fibred product $L\times_{X}\cdots\times_{X}L$ for the morphisms $t_{1}\circ\pi,\ldots,t_{s}\circ\pi:L\to X$. Equivalently, $F$ is the subvariety of $L^{s}$ consisting of elements $\ell\in L^{s}$ such that
$t_{1}(\pi(p_{1}(\ell)))=\cdots=t_{s}(\pi(p_{s}(\ell)))$
as elements of $X$. Applying the $G$-action to $\ell\in F$, we see that
$t_{i}(\pi(p_{i}(g\cdot\ell)))$ $=t_{i}h_{g,g^{-1}(i)}(\pi(p_{g^{-1}(i)}(\ell)))$
$=gt_{g^{-1}(i)}(\pi(p_{g^{-1}(i)}(\ell)))$
$=gt_{i}(\pi(p_{i}(\ell)))$
for every index $i$. This means that $g\cdot\ell\in L^{s}$ also represents an element from $F$ for every $g$. It also means that $G$ has an action on $F$ making the projection $F\to X$ equivariant. In other words, $\mathcal{F}$ has a $G$-linearization. ∎
The following theorem shows that, at least when $\operatorname{Pic}(X)$ is finitely generated, the numerical Amitsur group can actually be computed.
Theorem 5.4.
Suppose $\operatorname{Pic}(X)$ is finitely generated. Suppose $G$ is a finite group acting on $\operatorname{Pic}(X)$ such that $\chi$ is $G$-invariant. There exists a finite set $S\subset\operatorname{Pic}(X)$ such that
$\operatorname{Pic}^{\chi}(X,G)=\left\langle\chi(D)\Sigma_{G}(D)\ \middle|\ [D]\in S\right\rangle.$
Proof.
Let $H$ be a fixed subgroup of $G$ and consider
$Q_{H}:=\left\langle\chi(D)\Sigma_{G}(D)\ \middle|\ D\in\operatorname{Pic}(X)^{H}\right\rangle.$
The group $\operatorname{Pic}(X)^{H}$ is a finitely generated free abelian group and $\chi$ is a polynomial in the generators. Thus, by Proposition 3.3 there exists a finite set $S_{H}$ such that
$Q_{H}=\left\langle\chi(D)\Sigma_{G}(D)\ \middle|\ D\in S_{H}\right\rangle$
Now observe that
$\operatorname{Pic}^{\chi}(X,G)=\sum_{H\leq G}M_{H}$
where
$M_{H}:=\left\langle\chi(D)\Sigma_{G}(D)\ \middle|\ D\in\operatorname{Pic}(X),\ \operatorname{Stab}_{G}([D])=H\right\rangle.$
for each subgroup $H\leq G$. Note that $M_{H}\subseteq Q_{H}$, but we do not have equality in general.
If $[D]\in\operatorname{Pic}(X)$ and $H$ is a subgroup of $K=\operatorname{Stab}_{G}([D])$, then
$\sum_{g\in G/H}g[D]=[K:H]\sum_{g\in G/K}g[D]=[K:H]\Sigma_{G}(D).$
From this, we know that $Q_{H}\subseteq\sum_{H\leq K}M_{K}$ for all subgroups $H$ of $G$. Thus,
$\sum_{H\leq G}M_{H}=\sum_{H\leq G}Q_{H}$
and we conclude that
$\operatorname{Pic}^{\chi}(X,G)=\sum_{H\leq G}\left\langle\chi(D)\Sigma_{G}(D)\ \middle|\ D\in S_{H}\right\rangle.$
The set $S=\bigcup_{H\leq G}S_{H}$ is the desired set of generators. ∎
Remark 5.5.
It is interesting to consider which properties of $\operatorname{Am}(X,G)$ hold for $\operatorname{Am}^{\chi}(X,G)$ as well. In particular, the numerical Amitsur group is not a birational invariant even in the case where $G$ is trivial (consider $\mathbb{P}^{2}$ and $\mathbb{P}^{1}\times\mathbb{P}^{1}$). Additionally, it is not clear whether the canonical bundle $K_{X}$ is numerically split for every $X$, even though it is always linearizable.
6. Toric varieties
Let $X$ be a smooth projective toric variety [CLS11]. Suppose $T$ is the torus, $\Sigma$ is the fan, and $M:=\operatorname{Hom}(T,\mathbb{C}^{\times})$ is the character lattice. We have an exact sequence
(6.1) $0\to M\to\operatorname{TDiv}(X)\to\operatorname{Pic}(X)\to 0$
where $\operatorname{TDiv}(X)\cong\mathbb{Z}^{\Sigma(1)}\cong\mathbb{Z}^{r}$ is a free abelian group with basis indexed by the rays $\rho_{1},\ldots,\rho_{r}$ of $\Sigma$. Each ray corresponds to an irreducible $T$-invariant divisor of $X$. Let $D_{1},\ldots,D_{s}\in\operatorname{TDiv}(X)$ be a system of distinct representatives for the divisor classes in $\operatorname{Pic}(X)$ corresponding to these rays. For each $1\leq i\leq s$, let $n_{i}$ denote the number of rays in the class $[D_{i}]$.
From [Cox95], we recall the Cox ring and its connection to the automorphism group of $X$. The Cox ring $\operatorname{Cox}(X)$ of $X$ is a $\operatorname{Pic}(X)$-graded polynomial ring $\mathbb{C}[x_{1},\ldots,x_{r}]$ where each $x_{i}$ has the grading of the corresponding class in $\operatorname{Pic}(X)$. For every divisor class $[D]\in\operatorname{Pic}(X)$, we have
$H^{0}(X,\mathcal{O}_{X}(D))\cong\operatorname{Cox(X)}_{[D]}$
where $\operatorname{Cox(X)}_{[D]}$ denotes the homogeneous component of the Cox ring of degree $[D]$.
There is an exact sequence
$1\to S\to\widetilde{\operatorname{Aut}}(X)\to\operatorname{Aut}(X)\to 1$
where $S=\operatorname{Hom}(\operatorname{Pic}(X),\mathbb{C}^{\times})$ is the torus dual to the Picard group and $\widetilde{\operatorname{Aut}}(X)$ acts by polynomial automorphisms of $\operatorname{Cox}(X)$ that are compatible with the grading.
Lemma 6.1.
Suppose $G$ is a finite group acting on $X$ and let $\widetilde{G}$ be the pullback to $\widetilde{\operatorname{Aut}}(X)$. Let $\mathcal{L}$ be an effective line bundle in $\operatorname{Pic}(X)^{G}$ and let $\operatorname{ev}_{\mathcal{L}}:S\to\mathbb{C}^{\times}$ be the corresponding homomorphism. Then the extension defining the lifting group $G_{\mathcal{L}}$ is the image of the extension defining $\widetilde{G}$ along the induced morphism $(\operatorname{ev}_{\mathcal{L}})_{\ast}:H^{2}(G,S)\to H^{2}(G,\mathbb{C}^{\times})$.
Proof.
This amounts to proving that there is a commutative diagram with exact rows
[…]
Since $\mathcal{L}$ is $G$-invariant, the action of $\widetilde{G}$ leaves stable the homogeneous component $\operatorname{Cox}(X)_{[\mathcal{L}]}$. Since $\mathcal{L}$ is effective, $\operatorname{Cox}(X)_{[\mathcal{L}]}$ is non-zero and $\lambda\in S$ acts on $v\in\operatorname{Cox}(X)_{[\mathcal{L}]}$ via $\lambda\cdot v=\operatorname{ev}_{\mathcal{L}}(\lambda)v$. Since $\operatorname{Cox}(X)_{[\mathcal{L}]}\cong H^{0}(X,\mathcal{L})$, the result follows. ∎
From [Dun16, Theorem 4.5], the morphism $\alpha:\operatorname{Aut(X)}\to\operatorname{AutP}(X)$ has a splitting. Thus, we have the description
$\widetilde{\operatorname{Aut}}(X)\cong U\rtimes\left(\prod_{i=1}^{s}\operatorname{GL}_{n_{i}}(\mathbb{C})\right)\rtimes\operatorname{AutP}(X)$
where $U$ is a unipotent group. Each $\operatorname{GL}_{n_{i}}(\mathbb{C})$ acts on the linear span $V_{i}$ of the $n_{i}$ variables $\{x_{j}\}$ with degree $[D_{i}]$. The group $\operatorname{AutP}(X)$ acts by permuting the subspaces $\{V_{i}\}$. We have the description
$\operatorname{Aut}(\Sigma)=\left(\prod_{i=1}^{s}S_{n_{i}}\right)\rtimes\operatorname{AutP}(X)$
where each $S_{n_{i}}$ permutes the variables within each $V_{i}$. Taking the dual of (6.1) we obtain an exact sequence
$1\to S\to\widetilde{T}\to T\to 1$
of tori where $\widetilde{T}\cong\left(\mathbb{C}^{\times}\right)^{r}$ acts by scalar multiplication on each $x_{i}$.
Definition 6.2.
Suppose $J$ is a finite subgroup of $\operatorname{AutP}(X)$. Define
[…]
and define
$\operatorname{Am}^{T}(X,J):=\frac{\operatorname{Pic}(X)^{J}}{\operatorname{Pic}^{T}(X,J)}.$
Lemma 6.3.
Let $J$ be a subgroup of $\operatorname{AutP}(X)$ and let $W$ be the preimage of $J$ in $\operatorname{Aut}(\Sigma)$. There exist isomorphisms
(6.2) $H^{1}(W,M)\cong\operatorname{coker}\left(\operatorname{TDiv}(X)^{W}\to\operatorname{Pic}(X)^{J}\right)\cong\operatorname{Am}^{T}(X,J).$
Proof.
Observe that $\operatorname{TDiv}(X)$ is a permutation $W$-lattice; in other words, $W$ acts by permutations of a basis. By Shapiro’s Lemma, this means $H^{1}(W,\operatorname{TDiv}(X))=0$. We apply group cohomology $H^{i}(W,-)$ to (6.1) and obtain
$0\to M^{W}\to\operatorname{TDiv}(X)^{W}\to\operatorname{Pic}(X)^{J}\to H^{1}(W,M)\to 0$
This establishes the first isomorphism.
Let $H=\prod_{i=1}^{s}S_{n_{i}}$ denote the kernel of $W\to J$. The group $H$ acts on the set of rays $\{\rho_{i}\}$ by all permutations that leave invariant the divisor classes. Therefore, the $H$-orbit of $D_{i}$ is equal to the set of basis elements of $\operatorname{TDiv}(X)$ with degree $[D_{i}]$. Consequently,
$\beta_{1}:=\Sigma_{H}(D_{1}),\ \ldots,\ \beta_{s}:=\Sigma_{H}(D_{s})$
is a basis for $\operatorname{TDiv}(X)^{H}$. Therefore we have
$\operatorname{TDiv}(X)^{W}=\left(\operatorname{TDiv}(X)^{H}\right)^{J}$
where $J$ acts by permutations on $\beta_{1},\ldots,\beta_{s}$.
The divisor classes are all distinct, so the stabilizer of $J$ acting on $\beta_{i}$ is the same as that of $[D_{i}]$ in $\operatorname{Pic}(X)$. Therefore
$[\Sigma_{W}(D_{i})]=[\Sigma_{J}(\beta_{i})]=\Sigma_{J}([\beta_{i}])=n_{i}\Sigma_{J}([D_{i}]).$
We conclude that the image of $\operatorname{TDiv}(X)^{W}$ is equal to $\operatorname{Pic}^{T}(X,J)$. The second isomorphism now follows. ∎
Theorem 6.4.
For any finite subgroup $G$ of $\operatorname{Aut}(X)$ such that $\operatorname{AutP}(X,G)=J$, there exists a canonical surjection $\operatorname{Am}^{T}(X,J)\to\operatorname{Am}(X,G)$.
Proof.
Both groups are quotients of $\operatorname{Pic}(X)^{J}$ so it suffices to show that
$\operatorname{Pic}^{T}(X,J)\subseteq\overline{\operatorname{Pic}}(X,G).$
Thus, we only need to show that each $n_{i}\Sigma_{J}([D_{i}])$ is $G$-linearizable.
Let $\widetilde{G}$ be the preimage of $G$ in $\widetilde{\operatorname{Aut}}(X)$. Note that $\widetilde{G}$ is an extension of $G$ by the torus $S$. Since $G$ is finite, $\widetilde{G}$ is reductive. Thus, we may conjugate by an element of $U$ to ensure $\widetilde{G}$ is contained in
$R:=\left(\prod_{i=1}^{s}\operatorname{GL}_{n_{i}}(\mathbb{C})\right)\rtimes\operatorname{AutP}(X).$
(Alternatively, conjugation by an element of $U$ amounts to a different choice of torus $T$.)
Observe that $R$ acts on $\operatorname{Pic(X)}$ through $\operatorname{Aut}(X)$. Let $\widetilde{H}$ be the stabilizer in $\widetilde{G}$ of $[D_{i}]$ and let $H:=\widetilde{H}/S$ be its image in $\operatorname{Aut}(X)$. Since $R$ acts on the Cox ring, we see that $\widetilde{H}$ factors through the lifting group $H_{[D_{i}]}$ by Lemma 6.1. Observe that $\widetilde{H}$ leaves invariant the subspace $V_{i}$ of $\operatorname{Cox}(X)_{[D_{i}]}$ spanned by the variables $x_{j}$ with degree $[D_{i}]$. Therefore, there is a $n_{i}$-dimensional subrepresentation of the lifting group $H_{[D_{i}]}$ in $H^{0}(X,\mathcal{O}_{X}(D_{i}))$. Therefore, $n_{i}[D_{i}]$ is $H$-linearizable by Proposition 2.6.
By Lemma 5.3, we conclude that $n_{i}\Sigma_{G}([D_{i}])$ is $G$-linearizable. Since $J$ is simply the image of $G$, we have $n_{i}\Sigma_{G}([D_{i}])=n_{i}\Sigma_{J}([D_{i}])$ and the result is proven. ∎
Proposition 6.5.
For every subgroup $J$ of $\operatorname{AutP}(X)$, there exists a finite subgroup $G$ of $\operatorname{Aut}(X)$ such that $\operatorname{AutP}(X,G)=J$ and $\operatorname{Am}^{T}(X,J)\cong\operatorname{Am}(X,G)$.
Proof.
Let $W$ be the preimage of $J$ in $\operatorname{Aut}(\Sigma)$ and let $w=|W|$. Let $H$ be the $w$-torsion subgroup of $T$; thus, $H\cong(\mathbb{Z}/w\mathbb{Z})^{n}$. The group we will use is $G:=H\rtimes W$.
It suffices to show that
$\overline{\operatorname{Pic}}(X,G)\subseteq\operatorname{Pic}^{T}(X,J)$
as subgroups of $\operatorname{Pic}(X)^{J}$. Let $\mathcal{L}$ be a $G$-linearizable line bundle. We will prove that $\mathcal{L}$ is linearly equivalent to a divisor from $\operatorname{TDiv}(X)^{W}$, which will establish that $[\mathcal{L}]$ is in $\operatorname{Pic}^{T}(X,J)$ via Lemma 6.3.
Let $G_{\mathcal{L}}$ be the lifting group for $G$. Taking the preimage of $H$, we obtain the lifting group $H_{\mathcal{L}}$ as a subgroup of $G_{\mathcal{L}}$. We see that $G_{\mathcal{L}}/H_{\mathcal{L}}\cong G/H\cong W$. Let $\widetilde{G}$ be the pullback of $G$ to $\widetilde{\operatorname{Aut}}(X)$. Since $\operatorname{Aut}(\Sigma)$ acts on the Cox ring, the surjection $\widetilde{G}\to W$ has a splitting. Therefore, by Lemma 6.1, the surjection $G_{\mathcal{L}}\to W$ has a splitting. Thus $G_{\mathcal{L}}\cong H_{\mathcal{L}}\rtimes W$.
Let $T_{\mathcal{L}}$ be the lifting group of the torus $T$. Technically, we have only defined lifting groups for finite groups, but we only need it here to satisfy a version of Lemma 6.1. Specifically, we can just set $T_{\mathcal{L}}$ to be the quotient of $\widetilde{T}$ by the kernel of $\operatorname{ev}_{\mathcal{L}}:S\to\mathbb{C}^{\times}$. We have $(T\rtimes W)_{\mathcal{L}}\cong T_{\mathcal{L}}\rtimes W$ by similar reasoning as the previous paragraph.
We obtain a commutative diagram with exact rows:
[…]
In the above, the morphisms $H\rtimes W\to T\rtimes W$ and $H_{\mathcal{L}}\rtimes W\to T_{\mathcal{L}}\rtimes W$ are injective. The bottom row is the defining sequence for the lifting group $G_{\mathcal{L}}\cong H_{\mathcal{L}}\rtimes W$. Since $\mathcal{L}$ is $G$-linearizable, the bottom row is a split exact sequence.
Let $P:=\operatorname{Hom}(T_{\mathcal{L}},\mathbb{C}^{\times})$ and $Q:=\operatorname{Hom}(H_{\mathcal{L}},\mathbb{C}^{\times})$ be the character groups, with the induced $W$-action. Consider (6.3) where we only take the normal subgroup $N$ from each entry $N\rtimes W$, but remember that all the morphisms are $W$-equivariant. Next, we apply the character group functor $\operatorname{Hom}(-,\mathbb{C}^{\times})$ to the resulting diagram. We obtain a commutative diagram of $W$-modules with exact rows:
[…]
Exactness on the right follows in the first two rows since $\operatorname{Hom}(-,\mathbb{C}^{\times})$ is exact when restricted to tori. The bottom row is exact on the right since the morphism $P\to Q$ is surjective. Moreover, the bottom row is a split exact sequence of $W$-modules since the bottom row of (6.3) is a split exact sequence.
Recall that the group $\operatorname{Ext}^{1}_{\mathbb{Z}W}(\mathbb{Z},M)$ of extensions of $W$-modules is naturally isomorphic to $H^{1}(W,M)$. Let $\eta\in H^{1}(W,M)$ denote the extension defining $P$ in (6.4) and let $\nu\in H^{1}(W,M/wM)$ denote the extension defining $Q$. We have an exact sequence
[…]
of $W$-modules where $\nu=q_{\ast}(\eta)$. Recall from [Bro82, Corollary III.10.2] that multiplication by $w=|W|$ induces the zero map on group cohomology $H^{1}(W,-)$. Thus the induced long exact sequence
$\cdots\to H^{1}(W,M)\xrightarrow{-\times w}H^{1}(W,M)\xrightarrow{q_{\ast}}H^{1}(W,M/wM)\to\cdots$
establishes that $q_{\ast}$ is injective. Since the last row of (6.4) splits, we see that $\nu=0$ and, therefore $\eta=0$. Thus the middle row of (6.4) also splits.
Looking at the top right square of (6.4), we use the splitting to produce a $W$-equivariant composition
$\mathbb{Z}\to P\to\operatorname{TDiv}(X)\to\operatorname{Pic}(X)$
such that $1\mapsto[\mathcal{L}]$. Since $\mathbb{Z}$ has the trivial action, we conclude that $\mathcal{L}$ is in the image of $\operatorname{TDiv}(X)^{W}$ as desired. ∎
The groups $\operatorname{Am}^{T}(X,J)$ and $\operatorname{Am}^{\chi}(X,J)$ are both approximations for the usual Amitsur group that exploit Proposition 2.6 in similar ways. However, they are not the same in general (see Section 7.2 below). The basic reason for this is that the invariants $\dim H^{0}(X,D_{i})$, $\chi(D_{i})$ and $n_{i}$ are not equal in general. However, we have the following.
Proposition 6.6.
Let $X$ be a smooth Fano toric variety with reductive automorphism group. Then $\operatorname{Am}^{T}(X,J)\cong\operatorname{Am}^{\chi}(X,J)$ for every finite subgroup $J$ of $\operatorname{AutP}(X)$.
Proof.
Recall that the canonical bundle is
$K_{X}=-\sum_{i=1}^{r}D_{\rho_{i}}$
where $D_{\rho_{1}},\ldots,D_{\rho_{r}}$ are the divisors corresponding to the rays of $X$. Since $-K_{X}$ is ample, using a vanishing theorem of M. Mustaţă [CLS11, Theorem 9.3.7] we see that
$H^{p}\left(X,\mathcal{O}_{X}\left(D_{\rho_{j}}\right)\right)=H^{p}\left(X,\mathcal{O}_{X}\left(-K_{X}-\sum_{i\neq j}D_{\rho_{i}}\right)\right)=0$
for every index $1\leq j\leq r$ and every $p>0$. Therefore,
$\chi(D_{\rho_{j}})=\dim H^{0}(X,\mathcal{O}_{X}(D_{\rho_{j}}))$
for every divisor $D_{\rho_{j}}$ coming from a ray.
We recall the notion of Demazure roots in [Cox95, §4]. The unipotent radical of $\operatorname{Aut}(X)$ is nontrivial if and only if there exists a Demazure root that is not semisimple. Since $\operatorname{Aut}(X)$ is reductive, this means all the Demazure roots are semisimple. By [Cox95, Lemma 4.4], we conclude that the only monomials in $\operatorname{Cox}(X)$ with degree equal to that of the generators $x_{1},\ldots,x_{r}$ are the generators themselves. Therefore,
$H^{0}(X,\mathcal{O}_{X}(D_{i}))\cong\operatorname{Cox}(X)_{[D_{i}]}$
and we conclude that $\chi(D_{i})=n_{i}$ for all $1\leq i\leq s$.
Therefore, the defining generators of $\operatorname{Pic}^{T}(X,J)$ are contained in $\operatorname{Pic}^{\chi}(X,J)$ and we have a chain of surjections
$\operatorname{Am}^{T}(X,J)\to\operatorname{Am}^{\chi}(X,J)\to\operatorname{Am}(X,G)$
for every finite group $G$ with $\operatorname{AutP}(X,G)=J$. By Proposition 6.5, there exists a choice of $G$ such that they are in fact isomorphisms. ∎
Various conditions for a smooth Fano toric variety to have a reductive automorphism group can be found in [Nil06].
7. Examples
7.1. Products of projective spaces
Let $n$ and $m$ be positive integers and consider $X=(\mathbb{P}^{n})^{m}$. Viewing $X$ as a toric variety, one finds
$M\cong\mathbb{Z}^{nm},\ \operatorname{TDiv}(X)\cong\mathbb{Z}^{(n+1)m},\ \textrm{and }\operatorname{Pic}(X)\cong\mathbb{Z}^{m}.$
The basis of torus-invariant divisors are partitioned by linear equivalence into $m$ sets of size $n+1$. The full automorphism group is
$\operatorname{Aut}(X)\cong\operatorname{PGL}_{n+1}(\mathbb{C})\rtimes S_{m}$
while the automorphisms of the fan are given by
$\operatorname{Aut}(\Sigma)=(S_{n+1})^{m}\rtimes S_{m}.$
Thus, $\operatorname{AutP}(X)=S_{m}$. The automorphism groups are always reductive.
Suppose $J\subseteq S_{m}$. Then $W=(S_{n+1})^{m}\rtimes J$ and we find
$\operatorname{TDiv}(X)^{W}=(\mathbb{Z}^{(n+1)m})^{W}=\left(\left((\mathbb{Z}^{n+1})^{S_{n+1}}\right)^{m}\right)^{J}=\left(\mathbb{Z}^{m}\right)^{J}.$
The map into $\operatorname{Pic}(X)^{J}$ is simply multiplication by $n+1$. Thus, by Theorem 6.4, we find:
$\operatorname{Am}^{T}(X,J)\cong\operatorname{coker}\left(\operatorname{TDiv}(X)^{W}\to\operatorname{Pic}(X)^{J}\right)\cong\left((\mathbb{Z}/(n+1)\mathbb{Z})^{m}\right)^{J}.$
When $m=1$ we recover the fact that, for any finite group $G$ acting on $\mathbb{P}^{n}$, the group $\operatorname{Am}(\mathbb{P}^{n},G)$ is a cyclic group of order dividing $n+1$.
7.2. Hirzebruch surfaces
Let $X$ be the Hirzebruch surface of degree $e$. In other words, $X$ is a ruled surface over $\mathbb{P}^{1}$ with a section of self-intersection $-e$. Let $H$ be the class of the section and $F$ be the class of a fiber. In this case, $X$ is a toric surface with maximal rays
$D_{1}=\begin{pmatrix}1\\ 0\end{pmatrix},\ D_{2}=\begin{pmatrix}0\\ 1\end{pmatrix},\ D_{3}=\begin{pmatrix}-1\\ e\end{pmatrix},\ D_{4}=\begin{pmatrix}0\\ -1\end{pmatrix}$
where $[D_{1}]=F$, $[D_{2}]=H$, $[D_{3}]=F$, and $[D_{4}]=H+eF$. In this case, $\operatorname{Aut}(\Sigma)\cong S_{2}$ interchanges $D_{1}$ and $D_{3}$. Thus $\operatorname{AutP}(X)$ is trivial in this case. We compute that
$\operatorname{TDiv}(X)^{W}=\langle D_{1}+D_{3},D_{2},D_{4}\rangle,$
which has image $\langle 2F,H,H+eF\rangle$ in $\operatorname{Pic}(X)$. We conclude that
$\operatorname{Am}^{T}(X,G)\cong\begin{cases}0&\textrm{if $e$ is odd,}\\ \mathbb{Z}/2\mathbb{Z}&\textrm{if $e$ is even.}\end{cases}$
Hirzebruch surfaces are not necessarily Fano and their automorphism groups are typically not reductive. Thus, Theorem 6.6 does not apply and the numerical Amitsur group may have a different structure. Indeed, we see this in the following.
Proposition 7.1.
If $e$ is odd, then $\operatorname{Am}^{\chi}(X,G)$ is trivial. If $e$ is even, then $\operatorname{Am}^{\chi}(X,G)\cong(\mathbb{Z}/2)^{2}$.
Proof.
The class of the section $H$ and fiber $F$ form a basis for the Picard group with intersection theory $F^{2}=0$, $H^{2}=-e$, and $F\cdot H=1$. We have $K_{X}=-2H-(2+e)F$, so if $D=aH+bF$, then
$\chi(D)=-e\binom{a}{2}+ab+(1-e)a+b+1.$
We compute $\chi(D)D$ for $D\in\{H,F,H+F,H-F\}$ and determine that
$\begin{pmatrix}2-e\\ 0\end{pmatrix},\ \begin{pmatrix}0\\ 2\end{pmatrix},\ \begin{pmatrix}4-e\\ 4-e\end{pmatrix},\ \begin{pmatrix}-e\\ e\end{pmatrix}$
are all in $\operatorname{Pic}^{\chi}(X,G)$. If $e$ is odd, then these generate all of $\operatorname{Pic}(X)$ and therefore $\operatorname{Am}^{\chi}(X,G)$ is trivial.
If $e$ is even, then we conclude that $\langle 2H,2F\rangle$ is a subgroup of $\operatorname{Pic}^{\chi}(X,G)$. We will show that, in fact, equality holds. Let $D=aH+bF$ and consider $E:=\chi(D)D=cH+dF$. If $a$ is odd, then $a=2k+1$ and
$\chi(D)\equiv e(k+1)\equiv 0\mod{2},$
so $E$ has even coefficients. If $a$ is even, then
$\chi(D)\equiv-b+1\mod{2},$
which is even when $b$ is odd. Thus, if $e$ is even, then $c,d$ are even for all choices of $D$. Thus, $\operatorname{Pic}^{\chi}(X,G)$ never contains $H$, $H+F$ or $F$. ∎
7.3. The del Pezzo surface of degree 6
Here we determine the Amitsur groups of a del Pezzo surface $X$ of degree $6$. We recall that this is a toric variety obtained by blowing up $\mathbb{P}^{2}$ in three non-collinear points $p_{1},p_{2},p_{3}$. Let $E_{1},E_{2},E_{3}$ denote the exceptional divisors of the blown up points and let $L_{i}$ represent the strict transform of the line passing through the points $p_{j},p_{k}$ where $i,j,k\in\{1,2,3\}$ are distinct. The group $\operatorname{TDiv(X)}$ is the free abelian group with basis $\{E_{1},E_{2},E_{3},L_{1},L_{2},L_{3}\}$. All 6 divisors are in separate linear equivalence classes.
In this case, the fan $\Sigma$ in $\mathbb{R}^{2}$ is the unique complete fan with rays
$\begin{pmatrix}1\\ 0\end{pmatrix},\ \begin{pmatrix}0\\ 1\end{pmatrix},\ \begin{pmatrix}-1\\ -1\end{pmatrix},\ \begin{pmatrix}-1\\ 0\end{pmatrix},\ \begin{pmatrix}0\\ -1\end{pmatrix},\ \begin{pmatrix}1\\ 1\end{pmatrix},$
corresponding to the above basis for $\operatorname{TDiv(X)}$. The cocharacter lattice $N$ is isomorphic to $\mathbb{Z}^{2}$ and the matrices
$s=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\textrm{ and }r=\begin{pmatrix}0&1\\ -1&1\end{pmatrix}$
preserve the fan $\Sigma$. The fan has automorphism group $\operatorname{Aut}(\Sigma)=\langle s,r\rangle$ isomorphic to the dihedral group $D_{12}$ of order $12$, which can be thought of as the symmetries of the hexagon of exceptional lines. The full automorphism group is
$\operatorname{Aut}(X)\cong(\mathbb{C}^{\times})^{2}\rtimes D_{12},$
which coincides with the normalizer of the maximal torus in this case. In particular, $\operatorname{AutP}(X)\cong\operatorname{Aut}(\Sigma)$ and there is no distinction between $J$ and $W$ here.
Consider the exact sequence
$0\to M\to\operatorname{TDiv}(X)\to\operatorname{Pic}(X)\to 0$
where $M$ is the character lattice dual to $N$. The Picard group has rank $4$ and has basis $\{[H],[E_{1}],[E_{2}],[E_{3}]\}$ where $H$ is the strict transform of a general line on $\mathbb{P}^{2}$. We see that $L_{i}=H-E_{j}-E_{k}$ whenever $i,j,k$ are distinct.
By applying Theorem 6.4, the numerical Amitsur groups of $X$ can be computed. They are listed in Table 1.
$J$ Structure $\operatorname{Am}^{\chi}(X,J)$
$1$ $1$ 0
$\langle s\rangle$ $C_{2}$ 0
$\langle sr^{3}\rangle$ $C_{2}$ 0
$\langle r^{3}\rangle$ $C_{2}$ $(\mathbb{Z}/2\mathbb{Z})^{2}$
$\langle r^{2}\rangle$ $C_{3}$ $\mathbb{Z}/3\mathbb{Z}$
$\langle s,r^{3}\rangle$ $C_{2}^{2}$ $\mathbb{Z}/2\mathbb{Z}$
$\langle s,r^{2}\rangle$ $S_{3}$ 0
$\langle sr^{3},r^{2}\rangle$ $S_{3}$ $\mathbb{Z}/3\mathbb{Z}$
$\langle r\rangle$ $C_{6}$ 0
$\langle s,r\rangle$ $D_{12}$ 0
Table 1. Numerical Amitsur groups of $dP_{6}$
Remark 7.2.
Observe that both $\operatorname{Am}^{\chi}(X,D_{12})$ and $\operatorname{Am}^{\chi}(X,1)$ are trivial, but many of the intermediate subgroups have nontrivial Amitsur groups. This demonstrates that the assumption that $\operatorname{AutP}(X,G)=\operatorname{AutP}(X,H)$ cannot be removed in Proposition 2.12.
Remark 7.3.
In the cases where $\operatorname{Am}^{\chi}(X,J)\cong\mathbb{Z}/3\mathbb{Z}$, the exceptional divisors $E_{1},E_{2},E_{3}$ can be equivariantly blown down to produce $\mathbb{P}^{2}$. In the other cases where $\operatorname{Am}^{\chi}(X,J)$ is non-trivial, a pair of exceptional divisor $E_{i},L_{i}$ can be equivariantly blown down to produce $\mathbb{P}^{1}\times\mathbb{P}^{1}$. This indicates how one can find groups $G$ that realize non-trivial values for the ordinary Amitsur group $\operatorname{Am}(X,G)$ in these cases. Applying the contrapositive, we recover the fact that the Amitsur group is always trivial for a $G$-minimal del Pezzo $G$-surface of degree 6 (see [BCDP23, Proposition A.7]).
Remark 7.4.
It is tempting to try to compute the numerical Amitsur group via the quotient of the group
$P:=\langle\chi(\mathcal{L})\mathcal{L}\mid\mathcal{L}\in\operatorname{Pic}(X)^{G}\rangle\subseteq\operatorname{Pic}(X)^{G}$
instead of the definition of $\operatorname{Pic}^{\chi}(X,G)$ involving $G$-orbits from (5.2). However, they do not give the same result.
Every line bundle in $P$ is contained in $\operatorname{Pic}^{\chi}(X,G)$, but not conversely. Indeed, consider
$J=W=\langle sr^{3}\rangle=\left\langle\begin{pmatrix}0&-1\\ -1&0\end{pmatrix}\right\rangle\cong C_{2}.$
We see that
$\operatorname{TDiv}(X)^{W}=\langle E_{1}+L_{2},\ E_{2}+L_{1},\ E_{3}+L_{3}\rangle\cong\mathbb{Z}^{3},$
which maps onto
$\operatorname{Pic}(X)^{J}=\langle A:=H-E_{3},\ B:=H-E_{1}-E_{2}-E_{3}\rangle\cong\mathbb{Z}^{2}.$
Using Riemann-Roch and intersection theory, or using toric methods, we determine that
$\chi(mA+nB)=2mn-n^{2}+m+n+1.$
Thus, for $L=mA+nB$, we see that $\chi(L)[L]$ is given by
$\left(2m^{2}n-n^{2}m+m^{2}+mn+m\right)A+\left(2mn^{2}-n^{3}+mn+n^{2}+n\right)B.$
Observe that the coefficient of $A$ can be rewritten as
$2m^{2}n+m(n-n^{2})+(m^{2}+m),$
which is even for all values of $m,n$. Therefore $A$ is not in the span of $\chi(L)L$ for $L\in\operatorname{Pic}(X)^{G}$. However, $A=E_{1}+sr^{3}(E_{1})$ and $\chi(E_{1})=-1$. Thus $A\in\operatorname{Pic}^{\chi}(X,G)$.